If 1/x = 1/a + 1/b , determine the value of x. Hence prove sqrt(a-x/b-x) = a/b
Answer by rothauserc(4718)

1/x = 1/a + 1/b
:
1/x = (b+a)/ab
:
cross multiply fractions
:
x(b+a) = ab
:
************
x = ab/(b+a)
************
:
sqrt( (a-x)/(b-x) ) =
:
sqrt( (a-(ab/(b+a))/(b-(ab/(b+a)) ) =
:
sqrt( (ab+a^2-ab)/(b+a)/(b^2+ab-ab)/(b+a) ) =
:
sqrt( (a^2/(b+a)) / (b^2/(b+a)) ) =
:
sqrt( (a^2/(b+a)) * ((b+a)/b^2) ) =
:
sqrt( a^2/b^2) =
:
************************************
a/b or -a/-b = a/b, or -a/b or a/-b
************************************
: